Cole Company - company estimates that annual ufacturing overhead cost will be $600,000. Estimated annual operating ctivity ba please help
Answer #1

Solutions:

Cole Company

Overhead is applied on direct labor hours
Estimated Manufacturing Overhead cost 600000
Estimated Direct Labor Hours 40000
Predetermined Overhead Rate =Estimated Manufacturing Overhead Cost / Estimated Direct Labor Hours
=600000/40000
$15 per labor hour
Manufacturing Overhead applied =Actual Labor Hours * Predetermined Overhead Rate
=39800*15
$ 597000
Manufacturing Overhead Underapplied / (Overapplied) =Actual Overhead - Applied Overhead
=601000-597000
4000 Underapplied